OpenStudy (anonymous):

No idea how to do this... A rod has 3 notches in it as seen below (notches are where the numbers are). The rod is placed upon a fulcrum and your task is to place the 3 masses at the proper positions to achieve static equilibrium. The masses are: 20 grams, 17 grams, and 5 grams. <--4cm--><----7cm----><-----15cm----->

OpenStudy (anonymous):



OpenStudy (perl):

add the moments

OpenStudy (perl):

at the center of gravity, the moments add up to zero . A moment is equal to weight*displacement from center of gravity So we have -4*x + -7 * y + 15*z = 0 now x ,y,z are constrained 20,17,5 so plug these in and see if some combination equals to zero

OpenStudy (anonymous):

yup, 4cm – 20g 7cm – 5g 15cm – 17g

OpenStudy (perl):

wait i made a mistake

OpenStudy (perl):

we should have -(4+7)*x + -7 * y + 15*z = 0 -11x -7y + 15z = 0

OpenStudy (perl):

ok now I get x = 20 y = 5 z= 17

OpenStudy (perl):